An investment has a 20% chance of losing $20, a 30% chance of making $10, a 40% chance of making $20, and a 10% chance of making $100. What is the value of the investment?
Answer by VFBundy(438)

.20 * (-20) = -4
.30 * (+10) = +3
.40 * (+20) = +8
.10 * (+100) = +10

-4 + 3 + 8 + 10 = +17

The value of the investment is $17.